. If you wish to accumulate $197,000 in 5 years, how much must you deposit today in an account unding of interest?
Answer #1

Solution :- Time period = 5 * 12 / 6 = 10 Semi-annual periods. (Compounded semi-annually)

Rate of interest = 15 % * 6 / 12 = 7.50 % per half-year. (Compounded semi-annually)

Future value = Present value * (1 + Rate of interest)Time period

197000 = Present value * (1 + 7.50 / 100)10

197000 = Present value * (1 + 0.075)10

197000 = Present value * (1.075)10

197000 = Present value * 2.06103 (approx).

Present value = 197000 / 2.06103

Present value = $ 95583.28 (approx).

Conclusion :- Amount to be deposited today in account in the given question = $ 95583.28 (approx).